OpenStudy (anonymous):

if x^2+bx=(x+p)^2+q, express p and q in terms of b????

OpenStudy (campbell_st):

expand the right hand side and collect like terms \[x^2 + bx = x^2 + 2px + p^2 + q\] then \[bx = 2px + p^2 + q\] equating coefficients b = 2p and p^2 + q = 0

OpenStudy (anonymous):

my text gives the answer p=1/2b and q=-1/4b^2

OpenStudy (jamesj):

Are you sure you've given us the right equations then up top? Because what campbell has written is correct. (btw, what he has written also implies that q = -4b^2)

OpenStudy (jamesj):

e.g., is it x^2 + x/b = .... ?

OpenStudy (campbell_st):

the question is simply about equation coefficients of like terms...all you need to do is expand and compare...

OpenStudy (anonymous):

question 5

OpenStudy (asnaseer):

the answer that @campbell_st gave matches the answers given in your text book. just work through it, e.g.\[b=2p\implies p=\frac{1}{2}b\] now just re-arrange\[p^2+q=0\]

OpenStudy (campbell_st):

well is b = 2p then p = b/2 and is p^2 + q = 0 and p = B/2 then (b/2)^2 + q = 0 q =- b^2/4 just make p and q the subjects

OpenStudy (anonymous):

thank you guys.

OpenStudy (jamesj):

Oh right. I thought hc was saying p=1/2b and q=-1/4b^2, meaning p = 1/(2b), but no. In which case, asnaseer is exactly right.